# Structured Packing in LLM Training Improves Long Context Utilization

### v2

@@ Abstract
Recent advances in long-context Large Language Models (LCLMs) have generated significant interest, especially in applications such as querying scientific research papers. However, their potential is often limited by inadequate context utilization. We identify the absence of long-range semantic dependencies in typical training data as a primary hindrance. To address this, we delve into the benefits of frequently incorporating related documents into training inputs. Using the inherent directory structure of code data as a source of training examples, we demonstrate improvements in perplexity, even for tasks unrelated to coding. Building on these findings, but with a broader focus, we introduce Structured Packing for Long Context (SPLiCe). SPLiCe is an innovative method for creating training examples by using a retrieval method to collate the most mutually relevant documents into a single training context. Our results indicate that SPLiCe enhances model performance and can be used to train large models to utilize long contexts better. We validate our results by training a large $3$B model, showing both perplexity improvements and better long-context performance on downstream tasks.
@@ 1 Introduction
Large language models (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023; Bai et al., 2023) have completely transformed the field of AI and natural language processing. Long-context language models (LCLMs) are becoming increasingly popular and have the potential to unlock new capabilities (Anthropic, 2023). However, recent studies (Liu et al., 2023; Tworkowski et al., 2023) have highlighted numerous issues with their context utilization. Even if they have the technical ability to process long context, some models fail in practice, as evidenced by their poor performance even on synthetic retrieval tasks (Li et al., 2023a). Moreover, LLMs are often distracted from retrieving relevant information when provided with multiple documents in the prompt (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023), and typically struggle to utilize information from the middle of their input (Liu et al., 2023). There might be multiple reasons behind these issues, requiring structured research to address them.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components (e.g., the architecture and training objectives) unchanged. The fundamental question then is how to construct training examples such that using long context is beneficial for the model (e.g., it allows for obtaining lower perplexity or acquiring new downstream capabilities) while being non-trivial (like simple copying). We show that including multiple related documents in a training context can be beneficial - not only for evaluating on multiple documents, but also when considering language modeling perplexity on long, naturally occurring, single documents, which our evaluation is aimed for. First, we showcase this using the natural directory structure of the code data. Following that, we propose SPLiCe, Structured Packing for Long Context, that constructs training examples consisting of multiple similar documents selected using a retrieval method of choice (e.g., BM25).
@@
Figure 1: Training examples generated by the standard sampling procedure and proposed SPLiCe method.
We empirically validate SPLiCe on a large-scale code generation dataset (Li et al., 2023b), demonstrating improved perplexity evaluated at larger contexts, as well as a higher performance on TREC (Li & Roth, 2002; Hovy et al., 2001), Qasper, and synthetic task from (Liu et al., 2023). For that, we use the dataset generated by our method to continue pretraining an OpenLLaMA 3B LLM using the methodology described in (Tworkowski et al., 2023), resulting in a long-context language model with improved capabilities. Our contributions can be summarized as follows:
  • We show that structuring training data is a viable way of improving the context utilization of LLMs. We validate this using the natural directory structure of code data.
  • We introduce SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes retrieval (implemented using BM25 or Contriever (Izacard et al., 2022)) to find relevant documents without relying on directory structure. We show that using SPLiCe during training on code matches the model’s performance of the directory-structure method while being more generally applicable.
@@
  • We present a thorough analysis of the design choices of SPLiCe, the number of retrieved documents, and the order in which documents are merged into a training example.
  • We validate SPLiCe by training a large-scale model, OpenLLaMA 3B v2 (Geng & Liu, 2023). We show that SPLiCe leads to improved perplexity on long-context evaluation, improved in-context learning ability and better retrieval performance.
@@ 2 Method
SPLiCe is a method of constructing the training examples that allows improvement of large-scale models’ long-context utilization and, consequently, their performance on a range of tasks. Through experimental study, we positively validate that such an approach leads to better long-context capabilities.
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling, e.g. as demonstrated in the study by (Borgeaud et al., 2022) which highlights the advantages of incorporating global dependencies in retrieval-augmented language models. However, in typical datasets, the occurrence of such dependencies is rare (de Vries, 2023) and diminishes as the distance between elements increases, leading to challenges in model learning. Our approach addresses this challenge by focusing on increasing dependency density, achieved by creating training examples that consist of multiple related documents. The rationale of this strategy is reinforced by the work of Levine et al. (2022), which shows that the trained model establishes stronger dependencies between text segments present in the same training example.
In our work, we study the following methods:
Baseline
The standard approach, commonly used in LLM training pipelines (Brown et al., 2020), is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context (also known as example packing).
Repository-level code data
Certain datasets encompass meta-information regarding the data structure, which can be utilized to devise training examples. Take, for instance, a dataset consisting of GitHub repositories (e.g., (Li et al., 2023b)): rather than grouping documents randomly into a single context, we might group documents based on their subdirectory in a repository, thereby crafting a training example with potentially more cross-document dependencies compared to those generated randomly. This methodology, elaborated in (Wu et al., 2022), led to notable perplexity enhancements with augmented context length, suggesting that data orchestrated in such a manner could exhibit valuable long-context training signals. In our study, we strive to quantify the advantages of this methodology against the random packing baseline outlined above while maintaining the set of training documents to eliminate potential confounders.
Figure 2: Visualization of SPLiCe training example generation.
Structured Packing for Long Context (SPLiCe)
The additional structure provided by meta-information is unavailable in most datasets. We propose to overcome this challenge by using a retrieval method to identify related documents. We test two choices of retrieval methods BM25 (Robertson & Zaragoza, 2009; Bassani, 2023) and Contriever-MSMARCO (Izacard et al., 2022)). With those tools at our disposal, we propose SPLiCe, a method that creates training examples by building a tree of documents. It starts from a randomly sampled document and continues in a breadth-first manner, expanding each node with top-$k$ similar documents from the corpus. The final sequence consists of a flattened tree according to a certain traversal strategy (e.g., level-order traversal). An intuitive visualization of SPLiCe is presented in Figures 1 and 2, while a detailed pseudocode is given in Algorithm 1. We note that the general formulation of SPLiCe is flexible. For example, the parameter $k$ allows for interpolation between two modes of retrieval. For $k=1$ SPLiCe creates a path of related examples simulating a long document. For larger values of $k$ our method generates examples that are similar to the ones used by retrieval augmented models. Note that the dataset must be reasonably curated for SPLiCe to work efficiently. For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
@@
Input:
   $D$: the document corpus,
   $k$: the breadth hyper-parameter
   $L$: Maximum length of the returned training example
@@
$d_{r}\sim D$triangleright Sample the root document
@@
      $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$triangleright Retrieve top $k$ most similar documents to $d$ using a                                                         selected method, e.g., BM25
@@
         if $d_{i}\in D$ then $\triangleright$ RETRIEVE uses a precomputed index, and may return                                                     documents that are already in $C$
                  $C=C.\texttt{CONCAT}(d_{i})$triangleright Append $d_{i}$ to $C$
@@
return $\texttt{TRIM}(C,L)$
@@ 3 Experiments with medium-scale models
In this section, we conduct a comprehensive examination of the impact of document packing into training examples on long-context performance. We demonstrate that leveraging the inherent repository structure of code data can yield notable perplexity enhancements. Additionally, we establish that our proposed method, SPLiCe, either matches or surpasses this approach, exhibiting broader applicability to non-code data. We stress that we do not aim to surpass the approach to concatenate repository files together, instead we aim to show that SPLiCe is more general and can be used for non-code data. We carry out an ablation study on SPLiCe to scrutinize its design elements and hyperparameters in Appendix C.1. For this segment, due to computational constraints, we engage with $270$M models to ensure a meticulous investigation. In Section 4, we affirm that these findings extend to large-scale models.
@@ 3.1 Experimental setup
To investigate the impact of input data organization on enabling language models to leverage extended context, we employ $270$M models following a specific protocol. Initially, we pre-train a $270$M model on $6.3$B tokens from the original RedPajama mixture (TogetherComputer, 2023), using a context size of $2$K. Subsequently, we continue pretraining this model on $1$B tokens with an extended context length of $32$K, employing the Focused Transformer (FoT) objective Tworkowski et al. (2023), on a balanced mixture of the original RedPajama data (TogetherComputer, 2023) and long-context data assembled using the specified method. This approach is driven by practical factors: training with a shorter context expedites the process, and as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023), the model training can be continued on longer documents from the original data source while maintaining the desired long-context perplexity scaling. During the long-context pretraining phase, only the latter half of the mixture is altered, retaining the original RedPajama portion intact. This approach is loosely inspired by (Ouyang et al., 2022) and (Rozière et al., 2023) and aims to prevent the model from overfitting to artificially created long documents. Only the effect of data organization into training examples is tested, keeping other factors like architecture intact. For a detailed description of the model architecture, refer to Appendix A, and for information regarding the training data, see Appendix F.
@@
In the evaluation, we employ a context length of $32$K and measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets. The selection of these evaluation datasets is motivated by their inclusion of documents that stand to benefit from long-context information. Specifically, functions are often re-utilized, and similar terms are employed across papers. The advantage of long context in these datasets has been demonstrated for both arXiv (Chen et al., 2023) and StarCoder (Li et al., 2023b). We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. For further details regarding the chosen datasets and the evaluation setup, refer to Appendix F.
3.2 SPLiCe on code data
@@
In this section, we compare SPLiCe against the Baseline and repository-level code data (Repo) packing methods. We test two retrieval approaches for SPLiCe, the first one is BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), the second one is and Contriever-MSMARCO (Izacard et al., 2022)) and denote them as SPLiCe BM25 and SPLiCe Cont respectively. In SPLiCe BM25 retrieval takes into account the whole content of the document, whereas in SPLiCe Cont we use the first $512$ tokens and Faiss (Johnson et al., 2017b) for fast approximate inner-product search (for details see Appendix H). To make this comparison possible we use code data (required by Repo). We observe, see Table 1, that SPLiCe significantly outperforms the Baseline, and is either slightly better or on par with Repo. This is a promising result, as SPLiCe is more general and can be applied to non-code data (see Section 3.3).
Table 1: We train a $270$M parameter model on a 50/50 mixture of the RedPajama data (organized in a standard way) and code data (organized in one of four ways: SPLiCe BM25, SPLiCe Cont, Repo, Baseline – standard). We evaluate the perplexity of the models on arXiv and various subsets of StarCoder. The boldface denotes the best result within the training mixture. Note that SPLiCe is generally on par or better than Repo and beats the Baseline by a large margin. For detailed results, see Appendix B.
Training Method arXiv Code Code &
@@
C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
@@
Repo 5.47 3.22 2.83 2.24 2.96 3.12
C# SPLiCe BM25 5.52 3.33 2.90 2.46 3.11 3.26
SPLiCe Cont 5.53 3.35 2.91 2.48 3.12 3.27
@@
Repo 5.53 3.35 2.91 2.49 3.12 3.27
Python SPLiCe BM25 5.47 3.25 2.53 2.41 3.02 3.17
SPLiCe Cont 5.49 3.28 2.53 2.43 3.03 3.19
@@
Repo 5.48 3.27 2.54 2.44 3.03 3.18
To deepen the analysis, we prepared three subsets of the C code and ran the tested methods on each of them. In Table 2, we report the mean and standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results. For details about how the training sets were prepared please refer to Appendix F.
Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. For the C code we use three different subsets and report the mean and standard deviation.
Training Method arXiv Code Code &
@@
C SPLiCe BM25 5.463 $\pm$ 0.002 2.810 $\pm$ 0.002 2.942 $\pm$ 0.005 3.100 $\pm$ 0.004
SPLiCe Cont 5.477 $\pm$ 0.005 2.824 $\pm$0.001 2.957 $\pm$ 0.006 3.115 $\pm$ 0.006
Repo 5.474 $\pm$ 0.007 2.827 $\pm$ 0.006 2.958 $\pm$ 0.009 3.115 $\pm$ 0.009
3.3 SPLiCe on webtext data
The major advantage of our method over Repo is that SPLiCe can be applied to any textual data. Table 3 shows that using our method the model outperforms the baseline when trained on a $/50$ 50 mixture of RedPajama (packed in the standard way) and Wikipedia/StackExchange (prepared using SPLiCe). Notably, we observe that the model trained on StackExchange data significantly outperforms the one trained on C in the arXiv evaluation. Despite the non-code nature of StackExchange we still note perplexity improvements on code. Unsurprisingly, the perplexity on code is not as good as when training directly on code data. We conjecture that better data mixtures could bring further benefits.
Table 3: Perplexity results for training on a $/50$ 50 data mixture of RedPajama and one of {StackExchange,C} organized using mentioned methods. Note that when evaluating on arXiv the model trained with SPLiCe organized StackExchange outperforms the one trained on C.
Training Method arXiv Code Code &
data Haskell Python CUDA All arXiv
Wikipedia SPLiCe BM25 5.64 3.82 3.26 2.87 3.55 3.68
SPLiCe Cont 5.65 3.87 3.30 2.92 3.59 3.72
Baseline 5.73 3.97 3.37 3.00 3.68 3.81
StackExchange SPLiCe BM25 5.07 3.88 3.32 2.89 3.60 3.69
SPLiCe Cont 5.09 3.91 3.35 2.93 3.63 3.73
Baseline 5.14 3.94 3.36 2.94 3.65 3.74
C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
3.4 SPLiCe with different context extension approaches
Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension. FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. This allows for computationally efficient long-context fine-tuning. To show that our method works more generally, we have also evaluated it using more standard approaches. That is, we checked the effectiveness of the method when context extension is done in all layers, and parameters of Rotary Positional Encodings are either adjusted as in CodeLlama (Rozière et al., 2023), adjusted with YaRN (Peng et al., 2023), or left without changes. Table 4 shows the results.
Table 4: Perplexity results for training on a $/50$ 50 data mixture of RedPajama, and C# organized using the mentioned methods. We check how much SPLiCe can help when fine-tuning for the larger context using methods different than FoT (Tworkowski et al., 2023). That is, we test the effectiveness of SPLiCe when the context is extended to $16K$ in all layers and parameters of positional embedding are either adjusted as in CodeLlama (Rozière et al., 2023), with YaRN (Peng et al., 2023), or left without changes. The results indicate that SPLiCe still brings perplexity improvements in all of these setups. For YaRN we have used the parameters from (Peng et al., 2023) and adjusted the scale factor appropriately.
@@ 3.4 SPLiCe with different context extension approaches
Repo 6.25 4.87 3.56 2.85 3.74 3.89
CodeLlama C# SPLiCe BM25 5.74 4.28 3.22 2.53 3.34 3.49
@@
Repo 5.74 4.28 3.22 2.54 3.35 3.50
YaRN C# SPLiCe BM25 5.77 4.32 3.24 2.55 3.37 3.52
@@
Repo 5.77 4.32 3.24 2.56 3.38 3.53

### v5

@@ Abstract
Recent developments in long-context large language models have attracted considerable attention. Yet, their real-world applications are often hindered by ineffective context information use. This work shows that structuring training data to increase semantic interdependence is an effective strategy for optimizing context utilization. To this end, we introduce Structured Packing for Long Context (SPLiCe), a method for creating training examples by using information retrieval methods to collate mutually relevant documents into a single training context. We empirically validate SPLiCe on large $3$B and $7$B models, showing perplexity improvements and better long-context utilization on downstream tasks. Remarkably, already relatively short fine-tuning with SPLiCe is enough to attain these benefits. Additionally, the comprehensive study of SPLiCe reveals intriguing transfer effects such as training on code data leading to perplexity improvements on text data.
Machine Learning, Language Models, ICML
@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of Long-Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25 or repository structure) to collate mutually relevant documents into a single training context.
@@
Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe. Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2 and $7$Bv2 (Geng & Liu, 2023) for only $2$B–$6$B tokens already brings perplexity reduction. This reduction translates to substantial improvements in handling long-context information in downstream tasks that require retrieval and in-context learning. These tasks include TREC (Li & Roth, 2002; Hovy et al., 2001), DBpedia (Lehmann et al., 2015), Qasper (Shaham et al., 2022; Dasigi et al., 2021), HotPotQA (Yang et al., 2018), and the lost-in-the-middle benchmark (Liu et al., 2023). We perform a comprehensive study of the design choices and properties of SPLiCe, showing that the acquired long-context capabilities transfer between modalities, such as code and text.
Our contributions can be summarized as follows:
  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs. To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
  • We fine-tune OpenLLaMA $3$Bv2 and OpenLLaMA $7$Bv2 (Geng & Liu, 2023) using SPLiCe, showing that it improves perplexity on long context evaluation, in-context learning ability, and ability to retrieve relevant data from the context.
@@
  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
@@ 2 Method
SPLiCe is a method for constructing training examples that enhances the LLMs long-context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long context language modeling. We make the source code available to ensure the reproducibility of our results. (For this anonymous submission we share the zip file with our code.)
@@
  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training example   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., id, or random shuffle
@@
  $d_{r}\sim D$Sample the root document}
@@
     $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$
     {Retrieve top-$k$ most similar documents to $d$ using a selected method, e.g., BM25}
@@
        if $d_{i}\in D$ then
           {RETRIEVE uses a precomputed index, and may return documents that are already in $C$}
           $C=C.\texttt{APPEND}(d_{i})$Append $d_{i}$ to $C$}
@@
        end if
     end for
  end while
  return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from the fact that long-range dependencies are rare in training data (de Vries, 2023) and diminish with distance. It is thus unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. (Chan et al., 2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, (Shi et al., 2023b) shows that training on structured data improves the performance of LLMs on long-context tasks.
SPLiCe follows these intuitions. Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context. We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
Structured Packing for Long Context (SPLiCe)
We propose SPLiCe, a method that creates training examples by building a tree of documents, see Algorithm 1 and Figure 1 for a general overview.
SPLiCe starts by picking a random document from the dataset to create a single-node tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy.
The hyperparameter $k$ introduces flexibility and allows to interpolate between modes of retrieval. Indeed, for $k=1$ SPLiCe creates a path of related examples simulating a long document. For larger values of $k$, SPLiCe produces examples that are similar to the ones used by retrieval augmented models, e.g., (Borgeaud et al., 2022). Note that the dataset must be reasonably curated for SPLiCe to work efficiently. For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
  • SPLiCeRepo: based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together. A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
  • SPLiCeBM25: based on BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer model to rank documents based on their similarity to a query.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, the computational efficiency plays a non-trivial role. SPLiCeRepo is the fastest and easy to implement but requires additional directory structure, i.e., is not applicable to general web data. SPLiCeBM25 requires quadratic time index and needs dataset chunking to be applicable to large datasets, see Section 3.3. SPLiCeCont requires calculating embeddings for each document and retreval based on the cosine similarity. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Baseline The standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020).
Burstiness SPLiCe conjecturally falls into the framework presented in (Chan et al., 2022). This paper shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence . In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the baseline.
@@ 3 Experiments with Medium-Scale Models
In this section, we conduct a comprehensive examination of the impact of document packing on long-context performance, using medium-scale models (see Section 4 for large models results). We demonstrate that structuring data with SPLiCe is a generally applicable approach to improve long context performance. We study the performance in Section 3.2 and design choices of SPLiCe in Section 3.3.
@@ 3.1 Experimental Setup
In this part, we employ $270$M models and the following protocol. Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with an extended context length of $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. We employ the Focused Transformer (FoT) objective (Tworkowski et al., 2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. This aims to prevent the model from overfitting to artificially created long documents.
@@
In the evaluation, we measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by their inclusion of documents that can benefit from long-context information as demonstrated in (Chen et al., 2023) and (Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
We refer to Appendix A for a detailed description of the model architecture. For information regarding the training and evaluation data, see Appendix G.
3.2 Experimental Results
@@
In this section, we compare SPLiCe against Baseline. We test two retrieval approaches for SPLiCe: BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), and Contriever-MSMARCO (Izacard et al., 2022)) denoted as SPLiCeBM25 and SPLiCeCont respectively. In SPLiCeBM25 retrieval takes into account the whole content of the document, whereas in SPLiCeCont we use the first $512$ tokens and Faiss (Johnson et al., 2017) for fast approximate inner-product search (see Appendix H for details). By SPLiCeRepo, we denote the approach using the directory structure of code repositories. We also test our method with $64$K context length and report the results in Table 19 in Appendix J. In those experiments, we use $k=1$; see Section 3.3 for more details.
The results are presented in Table 3, Table 3, and Table 3, from which we draw the following conclusions.
Structuring data improves performance. We observe (see Table 3) that all variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe (see Table 3) positive effects when training on code and evaluating on arXiv and when training on web data and evaluating on code. This indicates that the positive effects of SPLiCe transfer between domains. We conjecture that better data mixtures could bring further benefits.
Retrieval method is of secondary importance. We observe (see Table 3) that the choice of retrieval method has a minor impact on the results. Typically, SPLiCeBM25 is slightly better than SPLiCeCont. On the code data, SPLiCeRepo is on par with SPLiCeBM25. We recall that SPLiCeRepo can be applied only to code data, whereas SPLiCeBM25 and SPLiCeCont are more general. In practice, we often observe that SPLiCeCont is faster than SPLiCeBM25 due to the use of Faiss.
[…]
Long Context Method arXiv Code Code &
@@
C SPLiCeBM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
SPLiCeCont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
SPLiCeRepo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
@@
C# SPLiCeBM25 5.52 (.13) 3.33 (.25) 2.90 (.17) 2.46 (.19) 3.11 (.20) 3.26 (.20)
SPLiCeCont 5.53(.12) 3.35(.23) 2.91(.16) 2.48(.17) 3.12(.19) 3.27(.19)
SPLiCeRepo 5.53 (.12) 3.35 (.23) 2.91 (.16) 2.49 (.16) 3.12 (.19) 3.27 (.19)
@@
Python SPLiCeBM25 5.47 (.10) 3.25 (.21) 2.53 (.09) 2.41 (.15) 3.02 (.15) 3.17 (.15)
SPLiCeCont 5.49 (.08) 3.28 (.18) 2.53 (.09) 2.43 (.13) 3.03 (.14) 3.19 (.13)
SPLiCeRepo 5.48 (.09) 3.27 (.19) 2.54 (.08) 2.44 (.12) 3.03 (.14) 3.18 (.14)
@@
Wikipedia SPLiCeBM25 5.64 (.09) 3.82 (.15) 3.26 (.11) 2.87 (.13) 3.55 (.13) 3.68 (.13)
SPLiCeCont 5.65 (.08) 3.87 (.10) 3.30 (.07) 2.92 (.08) 3.59 (.09) 3.72 (.09)
Baseline 5.73 3.97 3.37 3.00 3.68 3.81
StackExchange SPLiCeBM25 5.07 (.07) 3.88 (.06) 3.32 (.04) 2.89 (.05) 3.60 (.05) 3.69 (.05)
SPLiCeCont 5.09 (.05) 3.91 (.03) 3.35 (.01) 2.93 (.01) 3.63 (.02) 3.73 (.01)
Baseline 5.14 3.94 3.36 2.94 3.65 3.74
Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. We perform three trainings using different training datesets and report the mean and standard deviation.
Long Context Method arXiv Code Code &
@@
C SPLiCeBM25 5.463 $\pm$ .002 2.810 $\pm$ .002 2.942 $\pm$ .005 3.100 $\pm$ .004
SPLiCeCont 5.477 $\pm$ .005 2.824 $\pm$.001 2.957 $\pm$ .006 3.115 $\pm$ .006
SPLiCeRepo 5.474 $\pm$ .007 2.827 $\pm$ .006 2.958 $\pm$ .009 3.115 $\pm$ .009
Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
[…]
@@
CodeLlama C# SPLiCeBM25 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.53 (.05) 3.34 (.06) 3.49 (.06)
SPLiCeRepo 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.54 (.04) 3.35 (.05) 3.50 (.05)
@@
YaRN C# SPLiCeBM25 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.55 (.06) 3.37 (.07) 3.52 (.06)
SPLiCeRepo 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.56 (.05) 3.38 (.06) 3.53 (.05)
@@
SPLiCe works with various context extension results Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension. FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. This allows for computationally efficient long-context fine-tuning. To show that our method works more generally, we have also evaluated it using more standard approaches. That is, we checked the effectiveness of the method when context extension is done in all layers, and parameters of Rotary Positional Encodings are either adjusted as in CodeLlama (Rozière et al., 2023), adjusted with YaRN (Peng et al., 2023), or left without changes. Table 3 shows the results.
Training is stable. We prepared three subsets of the C code dataset and ran the tested methods on each of them. In Table 3, we report the mean perplexity and its standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
3.3 Design Choices Analysis
Choice of k As already mentioned in the previous section, the choice of the retrieval method (BM25 or Contriver) plays a minor role. In Appendix C.1, we study the role of $k$. We first observe that its choice has a minor impact on the results, with $k=1$ being the best choice. We leave for further work a more detailed analysis of this hyperparameter.
Document ordering We discovered that the ordering of the documents collected by the retrieval procedure might be a subtle issue. We consider three choices of ORDER function in Algorithm 1: identity (do not permute the documents), reverse (reverse the order of documents) and a random shuffle. In Appendix C.1, we present results for $270$M models showing negligible differences. However, we observed differences for large models, elaborated in Section 4.2.4, in which random shuffling is the best choice for the CodeLlama context extension method.
Corpus chunking Performing retrieval from the whole corpus might be cumbersome due to its scale. However, we find that chunking the corpus into smaller parts still brings improvements. In particular, for the experiments with English Wikipedia and BM25 presented in Table 3, we used a manageable size of $2$GB.